# Simulating the shaping of point-symmetric structures in the jittering jets explosion mechanism

### v2

@@ Abstract
We conduct three-dimensional hydrodynamical simulations of core-collapse supernovae by launching several pairs of jets into a collapsing core model and show that the jittering jets explosion mechanism (JJEM) can form a point-symmetric morphology that accounts for observed morphologies of many core-collapse supernovae (CCSN) remnants. Point-symmetric morphologies are composed of pairs of opposite structures around the center of the CCSN remnant. In the JJEM, the newly born neutron star launches several to a few tens of pairs of jets with stochastically varying directions, and these jets explode the star. In the simulations with the FLASH numerical code, we launch pairs of jets with varying directions, energies, opening angles, and durations into the massive stellar core and follow their evolution for about two seconds. We show that the jets form pairs of opposite filaments, clumps, bubbles, and lobes, namely, prominent point-symmetric morphologies. The interaction of the jets with the core leads to vigorous Rayleigh-Taylor instabilities and excites many vortices, which also shape clumps and filaments. Our results strengthen the claim that the JJEM is the primary explosion mechanism of CCSNe.
@@ 1 Introduction
The gravitational energy of the collapsing core powers core-collapse supernovae (CCSNe). Neutrinos carry most of the released gravitational energy as the neutron star (NS) cools down. A small fraction of this energy explodes the star. Two alternative theoretical models consider two different processes to deliver this energy to explode the star (for a recent reviews, see Soker 2024a; Janka 2025).
In the delayed neutrino explosion mechanism, a small fraction of the neutrinos interact with the collapsing core in the gain region, at $\approx 100{~{}\rm km}$ from the newly born NS, heat the material, revive the stalled shock and explode the star (e.g., Andresen et al. 2024; Boccioli & Fragione 2024; Burrows et al. 2024; Janka & Kresse 2024; Müller 2024; Müller et al. 2024; van Baal et al. 2024; Wang & Burrows 2024; Laplace et al. 2024; Huang et al. 2024; Boccioli et al. 2025b, a; Imasheva et al. 2025; Nakamura et al. 2025; Janka 2025; Wang & Burrows 2025, for some very recent studies). In the magnetorotational explosion mechanism, in rare cases of CCSNe, those that have rapidly rotating pre-collapse core (e.g., Shibagaki et al. 2024; Zha et al. 2024; Shibata et al. 2025, for recent studies of this mechanism), jets along a fixed axis explode the star; however, according to this mechanism, in most CCSNe neutrino heating explodes the star. Therefore, we consider the magnetorotational explosion mechanism part of the neutrino-driven mechanism, even though, for these rare cases, fixed-axis jets deliver most of the gravitational energy to the exploding star.
@@
One robust outcome of the JJEM is that many, but not all, CCSN remnants should possess point-symmetric morphology. Namely, the CCSN remnant has pairs of structural features opposite to the center of the CCSN remnant, and the pairs do not share their symmetry axes. The opposite structural features might be filaments, clumps, bubbles (faint zones closed by a brighter rim), lobes (bubbles with partial rims), and ears (protrusions from the main remnant’s shell having a decreasing cross-section with distance from the center). The neutrino-driven explosion mechanism cannot account for the point symmetric morphologies of so many CCSN remnants.
@@
The JJEM accounts for the point-symmetric morphology by pairs of jets that shape the supernova ejecta, hence the remnant. Many jets are choked inside the core, leaving no imprints on the remnants. Only one to several of the $N_{\rm 2j}\simeq 5-30$ pairs of exploding jets (Soker, 2024b) leave observable marks. In many cases, the point-symmetric morphology is not sharp and clear because several processes smear it (e.g., Soker & Shishkin 2025a), including interaction with circumstellar material, interaction with the interstellar medium, instabilities in the explosion process, the NS natal kick, heating of the ejecta by radioactive decay and the reverse shock, and the pulsar wind nebula. In recent years, studies have identified point-symmetric morphologies in over ten CCSN remnants and connected them to jet shaping, solidifying the JJEM as the primary and even sole explosion mechanism of CCSNe. Some of these point-symmetric CCSN remnants have a pair or more of opposite clumps, e.g., SNR 0540-69.3 (Soker, 2022b), Vela (Soker 2023; Soker & Shishkin 2025a), SN 1987A (Soker, 2024c, d), Cassiopeia A (Bear & Soker, 2025), W49B, (Soker & Shishkin, 2025b), and possibly also the Crab Nebula (Shishkin & Soker, 2024a). Other CCSN remnants do not show prominent pairs of clumps, these are N63A (Soker, 2024e), G321.3–3.9 (Soker, 2024f; Shishkin & Soker, 2024b), G107.7-5.1 (Soker, 2024f), the Cygnus Loop (Shishkin et al., 2024), W44 (Soker, 2024g), CTB 1 (Bear & Soker, 2023), and Puppis A (Bear et al., 2024). In attributing the point-symmetric pairs to shaping by jets, these studies applied a very common tool in the study of planetary nebulae, where studies consider jets to be behind the shaping of multipolar planetary nebulae and other point-symmetric morphologies (e.g., Morris 1987; Soker 1990; Sahai & Trauger 1998; Akashi & Soker 2018; Estrella-Trujillo et al. 2019; Tafoya et al. 2019; Balick et al. 2020; Rechy-García et al. 2020; Clairmont et al. 2022; Danehkar 2022; Moraga Baez et al. 2023; Ablimit 2024; Derlopa et al. 2024; Miranda et al. 2024; Sahai et al. 2024). The tool of identifying point-symmetric morphologies and connecting them to jet-driven explosions is new in the study of CCSNe. Our study is on that topic. In many planetary nebulae the jets precess, namely, continuously varying their direction (e.g., Guerrero & Manchado 1998; Miranda et al. 1998; Sahai et al. 2005; Boffin et al. 2012; Sowicka et al. 2017; Rechy-García et al. 2019; Guerrero et al. 2021; Clairmont et al. 2022). We postpone the study of precessing jets in CCSNe to a future study.
@@ 2.1 Grid and initial conditions
We utilize a fully three-dimensional AMR grid with 8 refinement levels, resulting in $2^{8}$ cells along each axis. The computational domain is a Cartesian grid $(x,y,z)$ with outflow boundary conditions applied at all boundaries. The plane $z=0$ is set as the equatorial plane of both the star and the surrounding medium. Our simulations cover the entire spatial domain, including both sides of the equatorial plane.
@@
We include the gravity of a point mass of $1.4M_{\odot}$ at the center to represent the newly born NS. This gravitational potential influences the surrounding gas, particularly affecting the backflow. For numerical reasons that allow us to conduct six simulations, we inject the jets at $500{~{}\rm km}$ and not at $50{~{}\rm km}$. We also do not include NS wind nor later jets that result from the accretion of the back-flowing material. For these reasons, in this first study in a series, we do not analyze the flow near the center, namely inner $\simeq 15\%$ of the grid size, which is the volume inner to $\simeq 5000{~{}\rm km}$.
@@ 2.3 Simulated cases
Simulation A3 consists of three pairs of jets, forming a pyramid shape, and launched sequentially. Table 1 lists the parameters of the pairs. The jets have the same half-opening angle $\alpha_{\rm j}$, velocity $v_{\rm j}$, and mass loss rate $\dot{m}_{\rm j}$. Note that the last pair lasts longer than the other pairs. The total kinetic energy carried by the six jets in this simulation is $E_{\rm k,tot}=2.5\times 10^{51}\ \rm erg$ (the thermal energy adds $20\%$ of the kinetic energy). Simulation A5 consists of five pairs of jets, with their properties listed in Table 2. The first three pairs in this simulation are identical to those in simulation A3. There is a quiet period between consecutive jet-launching episodes. Note that the last pair lasts longer and has a larger half-opening angle than the other pairs. The total kinetic energy carried by the ten jets in this simulation is $E_{\rm k,tot}=1.5\times 10^{51}\ \rm erg$.
@@
Notes: The initial parameters of simulation A3 of three pairs of jets. In all simulations, the launching velocity of the jets is $v_{\rm j}=50,000{~{}\rm km}{~{}\rm s}^{-1}$. The second and third rows give each jet-launching episode’s initial and final time. $(\theta_{\rm j},\phi_{\rm j})$ is the direction of the symmetry axis of the pair, where $\theta_{\rm j}$ is the polar angle, i.e., the angle to the $z$ axis in our grid, and $\phi_{\rm j}$ is the azimuthal angle, i.e., the angle of the projected axis on the $xy$ plane relative to the $x$ axis; $\alpha_{\rm j}$ is the half opening angle of the jets; $\dot{m}_{\rm j}$ is the mass loss rate of one jet; $E_{\rm k,2j}$ is the kinetic energy of the pair of jets; thermal energy is $20\%$ of that. Note that the last pair lasts longer than the other pairs. The total kinetic energy carried by the jets in this simulation is $E_{\rm k,tot}({\rm A3})=2.5\times 10^{51}\ \rm erg$.
@@ 3.1 Density structures
Figure 1 presents the density of simulation A3 (see Table 1) in three planes, from top to bottom $z=0$, $x=0$ and $y=0$, and at three times as indicated in the upper row. Figure 2 presents the density maps of simulation A5; note the different density scales of the color bars and the different times of the maps. The total energy in simulation A3 is larger than in simulation A5, hence it evolves faster.
@@
Figures 1 and 2 present the point-symmetric density structure that the jets shape in the core. Papish & Soker (2014) already obtained similar results, but we use a larger grid and more cases. The 2D density maps in the planes show bubbles (low-density zones closed with a high-density rim), high-density elongated structures, i.e., filaments, and high-density clumps, in a point symmetric morphology. Namely, each structural feature has a partner on the other side of the center. The last jet activity in simulation A3 was at $t_{\rm f,j}=0.25{~{}\rm s}$ and in simulation A5 at $t_{\rm f,j}=0.413{~{}\rm s}$. Figures 1 and 2 show that the point-symmetric morphologies in the two simulations remain prominent for five times as long. The point-symmetric structure reaches distances from the center of $\simeq 20,000-30,000{~{}\rm km}$. Beyond those distances, the density sharply drops and we expect the point-symmetric structures to stay in the CCSN remnant phase. Many processes will smear the point-symmetric structure, including the NS kick, instabilities, interaction with the circumstellar material, and heat release by radioactive decay that forms bubbles (nickel bubbles). Therefore, identifying point-symmetric morphologies in CCSN remnants might be challenging (see references in Section 1).
To further illustrate the complex point-symmetric morphologies, in Figures 3 - 5 we present 3D maps of the six simulations from three different points of view. Each 3D map presents three equidensity surfaces, in yellow (low density), red, and blue (high density; see captions). The colored surfaces are partially transparent, so one can see deeper surfaces. Because the colored-surfaces are not fully transparent, zones away from the observer appear differently than those closer to the observer. Therefore, in some cases there is no full symmetry between opposite sides in those maps.
@@
The 3D density surface maps, Figures 3 - 5, clearly present the point-symmetric morphologies by emphasizing pairs of structural features, particularly bubbles and clumps. Bubbles are prominent, as example, in the lower right panel of Figure 3 for simulation B7 (two opposite white areas), and in the three lower panels of Figure 5, simulations A5, B5, and B7 (the pairs of yellow areas surrounded by red rims). The upper left panel of Figures 3 - 5 show the large bubbles of simulation A3 that we present in Figure 1.
Pairs of clumps are prominent in simulations P1 and P2 that we present in the upper middle and upper right panels of Figures 3 - 5, particularly in Figure 5. Inner pairs of clumps are also prominent, as small blue areas in the larger red zones, in simulations A5 and B5, as seen in the lower left and middle panels in Figures 3 - 5. The upper middle panel of Figure 5 is for simulation P1 with three pairs of jets. However, we see five pairs of clumps, and each of the 10 clumps is fragmented to smaller knots, due to instabilities that we discuss in Section 4. The finding that three pairs of jets can form five pairs of clumps is highly significant when analyzing observations. Soker & Shishkin (2025a) who identified seven pairs of clumps in the Vela CCSN remnant, already pointed out that two pairs result from one pair of jets. Bear & Soker (2025) who identified the point symmetric structure of Cassiopeia A with nine pairs of structural features, proposed that a pair of jets can form two or more pairs of clumps and filaments. Our results solidify the claim from observations that a pair of jets can form two or more pairs of clumps and filaments. Or, more accurately, the interaction between two pairs of jets can form more than two pairs of clumps.

### v3

@@ Abstract
We conduct three-dimensional hydrodynamical simulations of core-collapse supernovae by launching several pairs of jets into a collapsing core model and show that the jittering jets explosion mechanism (JJEM) can form a point-symmetric morphology that accounts for observed morphologies of about a dozen core-collapse supernovae (CCSN) remnants. Point-symmetric morphologies are composed of pairs of opposite structures around the center of the CCSN remnant. In the JJEM, the newly born neutron star launches several to a few tens of pairs of jets with stochastically varying directions, and these jets explode the star. In the simulations with the FLASH numerical code, we launch pairs of jets with varying directions, energies, opening angles, and durations into the massive stellar core and follow their evolution for about two seconds. We show that the jets form pairs of opposite filaments, clumps, bubbles, and lobes, namely, prominent point-symmetric morphologies. The interaction of the jets with the core leads to vigorous Rayleigh-Taylor instabilities and excites many vortices, which also shape clumps and filaments. Our results suggest that the JJEM could play a central role in the explosion mechanism of CCSNe; neutrino heating can boost the role of jets.
@@ 1 Introduction
The gravitational energy of the collapsing core powers core-collapse supernovae (CCSNe). Neutrinos carry most of the released gravitational energy as the neutron star (NS) cools down. A small fraction of this energy explodes the star. Two alternative theoretical models consider two different processes to deliver this energy to explode the star (for recent reviews, see Soker 2024a; Janka 2025).
In the delayed neutrino explosion mechanism, a small fraction of the neutrinos interact with the collapsing core in the gain region, at $\approx 100{~{}\rm km}$ from the newly born NS, heat the material, revive the stalled shock and explode the star (e.g., Andresen et al. 2024; Boccioli & Fragione 2024; Burrows et al. 2024; Janka & Kresse 2024; Müller 2024; Müller et al. 2024; van Baal et al. 2024; Wang & Burrows 2024; Laplace et al. 2024; Huang et al. 2024; Boccioli et al. 2025b, a; Imasheva et al. 2025; Nakamura et al. 2025; Janka 2025; Wang & Burrows 2025, for some very recent studies). In the magnetorotational explosion mechanism, in rare cases of CCSNe, those that have a rapidly rotating pre-collapse core (e.g., Shibagaki et al. 2024; Zha et al. 2024; Shibata et al. 2025, for recent studies of this mechanism), jets along a fixed axis explode the star; however, according to this mechanism, in most CCSNe, neutrino heating explodes the star. Therefore, we consider the magnetorotational explosion mechanism part of the neutrino-driven mechanism, even though, for these rare cases, fixed-axis jets deliver most of the gravitational energy to the exploding star.
@@
One robust outcome of the JJEM is that many, but not all, CCSN remnants should possess point-symmetric morphology. Namely, the CCSN remnant has pairs of structural features opposite to the center of the CCSN remnant, and the pairs do not share their symmetry axes. The opposite structural features might be filaments, clumps, bubbles (faint zones closed by a brighter rim), lobes (bubbles with partial rims), and ears (protrusions from the main remnant’s shell having a decreasing cross-section with distance from the center). The neutrino-driven explosion mechanism cannot account for the point-symmetric morphologies of about a dozen CCSN remnants.
@@
The JJEM accounts for the point-symmetric morphology by pairs of jets that shape the supernova ejecta, hence the remnant. Many jets are choked inside the core, leaving no imprints on the remnants. Only one to several of the $N_{\rm 2j}\simeq 5-30$ pairs of exploding jets (Soker, 2024b) leave observable marks. In many cases, the point-symmetric morphology is not sharp and clear because several processes smear it (e.g., Soker & Shishkin 2025a), including interaction with circumstellar material, interaction with the interstellar medium, instabilities in the explosion process, the NS natal kick, heating of the ejecta by radioactive decay and the reverse shock, and the pulsar wind nebula. In recent years, studies have identified point-symmetric morphologies in over ten CCSN remnants and connected them to jet shaping, solidifying the JJEM as the primary and even sole explosion mechanism of CCSNe. Some of these point-symmetric CCSN remnants have a pair or more of opposite clumps, e.g., SNR 0540-69.3 (Soker, 2022b), Vela (Soker 2023; Soker & Shishkin 2025a), SN 1987A (Soker, 2024c, d), Cassiopeia A (Bear & Soker, 2025), W49B, (Soker & Shishkin, 2025b), and possibly also the Crab Nebula (Shishkin & Soker, 2024a). Other CCSN remnants do not show prominent pairs of clumps; these are N63A (Soker, 2024e), G321.3–3.9 (Soker, 2024f; Shishkin & Soker, 2024b), G107.7-5.1 (Soker, 2024f), the Cygnus Loop (Shishkin et al., 2024), W44 (Soker, 2024g), CTB 1 (Bear & Soker, 2023), Puppis A (Bear et al., 2024), and SNR G0.9+0.1 (Soker, 2025). In attributing the point-symmetric pairs to shaping by jets, these studies applied a very common tool in the study of planetary nebulae, where studies consider jets to be behind the shaping of multipolar planetary nebulae and other point-symmetric morphologies (e.g., Morris 1987; Soker 1990; Sahai & Trauger 1998; Akashi & Soker 2018; Estrella-Trujillo et al. 2019; Tafoya et al. 2019; Balick et al. 2020; Rechy-García et al. 2020; Clairmont et al. 2022; Danehkar 2022; Moraga Baez et al. 2023; Ablimit 2024; Derlopa et al. 2024; Miranda et al. 2024; Sahai et al. 2024). The tool of identifying point-symmetric morphologies and connecting them to jet-driven explosions is new in the study of CCSNe. Our study is on that topic. In many planetary nebulae, the jets precess, namely, continuously varying their direction (e.g., Guerrero & Manchado 1998; Miranda et al. 1998; Sahai et al. 2005; Boffin et al. 2012; Sowicka et al. 2017; Rechy-García et al. 2019; Guerrero et al. 2021; Clairmont et al. 2022). We postpone the study of precessing jets in CCSNe to a future study.
@@ 2.1 Grid and initial conditions
We utilize a fully three-dimensional AMR grid with 8 refinement levels, resulting in $2^{10}$ cells along each axis. The computational domain is a Cartesian grid $(x,y,z)$ with outflow boundary conditions applied at all boundaries. The plane $z=0$ is set as the equatorial plane of both the star and the surrounding medium. Our simulations cover the entire spatial domain, including both sides of the equatorial plane.
@@
We include the gravity of a point mass of $1.4M_{\odot}$ at the center to represent the newly born NS. This gravitational potential influences the surrounding gas, particularly affecting the backflow. For numerical reasons that allow us to conduct six simulations, we inject the jets at $500{~{}\rm km}$ and not at $50{~{}\rm km}$. We also do not include NS wind or later jets that result from the accretion of the back-flowing material. For these reasons, in this first study in a series, we do not analyze the flow near the center, namely the inner $\simeq 15\%$ of the grid size, which corresponds to the volume inner to $\simeq 5000{~{}\rm km}$. According to the JJEM, the accretion disk (or belt) launches the jets at $20-50{~{}\rm km}$ from the center. The jets interact with the inflowing core material both below and above the stalled shock, in the region of $\simeq 50-200{~{}\rm km}$. This turbulent inflow, along with the unrelaxed intermittent accretion disk that launches the jets, leads to the two opposite jets of each jet-launching episode not being equal when they emerge at hundreds of kilometers from the center. Other than that, there are no large effects on the role of the jets in exploding the core. Therefore, injecting the jets at $500{~{}\rm km}$ is adequate for this study. In this study, we inject equal jets in each pair. In future simulations, we will inject two unequal opposite jets to better reflect the expected behavior according to the JJEM.
@@ 2.3 Simulated cases
Simulation A3 consists of three pairs of jets launched sequentially, forming a pyramid shape. Table 1 lists the parameters of the pairs. The jets have the same half-opening angle $\alpha_{\rm j}$, velocity $v_{\rm j}$, and mass loss rate $\dot{m}_{\rm j}$. Note that the last pair lasts longer than the other pairs. The total kinetic energy carried by the six jets in this simulation is $E_{\rm k,tot}=2.5\times 10^{51}\ \rm erg$ (the thermal energy adds $20\%$ of the kinetic energy). Simulation A5 consists of five pairs of jets, with their properties listed in Table 2. The first three pairs in this simulation are identical to those in simulation A3. There is a quiet period between consecutive jet-launching episodes. Note that the last pair lasts longer and has a larger half-opening angle than the other pairs. The total kinetic energy carried by the ten jets in this simulation is $E_{\rm k,tot}=1.5\times 10^{51}\ \rm erg$.
@@
Notes: The initial parameters of simulation A3, which has three pairs of jets. In all simulations, the launching velocity of the jets is $v_{\rm j}=50,000{~{}\rm km}{~{}\rm s}^{-1}$. The second and third rows provide the initial and final times for each jet-launching episode. $(\theta_{\rm j},\phi_{\rm j})$ is the direction of the symmetry axis of the pair, where $\theta_{\rm j}$ is the polar angle, i.e., the angle to the $z$ axis in our grid, and $\phi_{\rm j}$ is the azimuthal angle, i.e., the angle of the projected axis on the $xy$ plane relative to the $x$ axis; $\alpha_{\rm j}$ is the half opening angle of the jets; $\dot{m}_{\rm j}$ is the mass loss rate of one jet; $E_{\rm k,2j}$ is the kinetic energy of the pair of jets; thermal energy is $20\%$ of that. Note that the last pair lasts longer than the other pairs. The total kinetic energy carried by the jets in this simulation is $E_{\rm k,tot}({\rm A3})=2.5\times 10^{51}\ \rm erg$.
@@ 3.1 Density structures
Figure 1 presents the density of simulation A3 (see Table 1) in three planes, from top to bottom $z=0$, $x=0$, and $y=0$, and at three times as indicated in the upper row. Figure 2 presents the density maps of simulation A5; note the different density scales of the color bars and the different times of the maps. The total energy in simulation A3 is larger than in simulation A5, hence it evolves faster.
@@
Figures 1 and 2 present the point-symmetric density structure that the jets shape in the core. Papish & Soker (2014) already obtained similar results, but we use a larger grid and more cases. The 2D density maps in the planes show bubbles (low-density zones closed with a high-density rim), high-density elongated structures, i.e., filaments, and high-density clumps, in a point-symmetric morphology. Namely, each structural feature has a partner on the other side of the center. The last jet activity in simulation A3 was at $t_{\rm f,j}=0.25{~{}\rm s}$ and in simulation A5 at $t_{\rm f,j}=0.413{~{}\rm s}$. Figures 1 and 2 show that the point-symmetric morphologies in the two simulations remain prominent for five times as long. The point-symmetric structure reaches distances from the center of $\simeq 20,000-30,000{~{}\rm km}$. Beyond those distances, the density sharply drops and we expect the point-symmetric structures to stay in the CCSN remnant phase. Many processes will smear the point-symmetric structure, including the NS kick, instabilities, interaction with the circumstellar material, and heat release by radioactive decay that forms bubbles (nickel bubbles). Therefore, identifying point-symmetric morphologies in CCSN remnants might be challenging (see references in Section 1).
To further illustrate the complex point-symmetric morphologies, in Figures 3 - 5 we present 3D maps of the six simulations from three different points of view. Each 3D map presents three equidensity surfaces, in yellow (low density), red, and blue (high density; see captions). The colored surfaces are partially transparent, so one can see deeper surfaces. Because the colored-surfaces are not fully transparent, zones away from the observer appear differently than those closer to the observer. Therefore, in some cases, there is no full symmetry between opposite sides in those maps.
@@
The 3D density surface maps, Figures 3 - 5, clearly present the point-symmetric morphologies by emphasizing pairs of structural features, particularly bubbles and clumps. Bubbles are prominent, as example, in the lower right panel of Figure 3 for simulation B7 (two opposite white areas), and in the three lower panels of Figure 5, simulations A5, B5, and B7 (the pairs of yellow areas surrounded by red rims). The upper left panels of Figures 3 - 5 show the large bubbles of simulation A3 that we present in Figure 1.
Pairs of clumps are prominent in simulations P1 and P2 that we present in the upper middle and upper right panels of Figures 3 - 5, particularly in Figure 5. Inner pairs of clumps are also prominent, as small blue areas in the larger red zones, in simulations A5 and B5, as seen in the lower left and middle panels in Figures 3 - 5. The upper middle panel of Figure 5 is for simulation P1 with three pairs of jets. However, we see five pairs of clumps, and each of the 10 clumps is fragmented to smaller knots, due to instabilities that we discuss in Section 4. The finding that three pairs of jets can form five pairs of clumps is highly significant when analyzing observations. Soker & Shishkin (2025a) who identified seven pairs of clumps in the Vela CCSN remnant, already pointed out that two pairs result from one pair of jets. Bear & Soker (2025) who identified the point-symmetric structure of Cassiopeia A with nine pairs of structural features, proposed that a pair of jets can form two or more pairs of clumps and filaments. Our results solidify the claim from observations that a pair of jets can form two or more pairs of clumps and filaments. Or, more accurately, the interaction between two pairs of jets can form more than two pairs of clumps.